OpenStudy (anonymous):

Someone please help me? Solve: 4(x + 1) = −4

OpenStudy (anonymous):

use that property a(b+c)=ab+ac what will you get?

OpenStudy (anonymous):

@Micheellexox ??

OpenStudy (anonymous):

can't you write 4(x+1) as 4*x + 4*1 by using the property a(b+c)=a*b+a*c